OpenStudy (anonymous):

What are the x- and y-intercepts of the graph of y = x2 + 3x - 10? A) x-intercepts: (-2, 0) and (5, 0) y-intercept: (0, 10) B) x-intercepts: (-2, 0) and (-5, 0) y-intercept: (0, 10) C) x-intercepts: (-2, 0) and (5, 0) y-intercept: (0, -10) D) x-intercepts: (2, 0) and (-5, 0) y-intercept: (0, -10)

OpenStudy (anonymous):

set x=0 for the y-intercept: y = 0^2 + 3*0 -10 set y = 0 for the x-intercept: 0 = x^2 + 3*x -10... solve for x. you might want to factor that quadratic to easily find the x-intercepts.